# CLMFormer: Mitigating Data Redundancy to Revitalize Transformer-based Long-Term Time Series Forecasting System

### v4

@@ III-A1 Input Embedding
The input data are first passed through an embedding. Different from other types of sequential data, in time series problems, the model should capture the time information. This is extremely crucial in LSTF settings, as models are required to make predictions up to weeks in advance. We decided to utilize the time-series-specific embedding method proposed by [15]. The method applies a combination of the context vector, the positional embedding, and the seasonal embedding. The context vector $u^{t}\in R^{L_{S}*d_{model}}$ is obtained by projecting the model input $X^{t}\in R^{L_{S}*d_{f}}$ into a dimension of $d_{model}$ by an 1-D convolutional layer. The goal of the context vector is to transform the input data into an appropriate dimension. The positional embedding (PE(pos)) and the seasonal embedding (SE(pos)) are used to capture the local context and the global time feature, respectively. The formula for the input embedding is:
$X_{embed}^{t}=\delta u^{t}+\rm{PE}(\rm{pos})+\sum_{k}[\rm{SE(pos)}]_{k}$ (2)
@@
III-A2 Encoder-Decoder Architecture
The backbone of our method lies in Transformer-based models following an encoder-decoder structure. Despite some minor differences between different variants, their general purpose is the same. The encoder takes past input sequences into the model and generates hidden states as outputs that contain time pattern information, which the decoder uses to make predictions. We decided to use generative-style decoding [15], which has been proven to achieve superior performances in LSTF problems and have exponentially better computational efficiency than conventional dynamic decoding [34]. The generative-style decoding generates all predictions in one step, with the decoder feeding vector $X^{t}_{feed,dec}$ presented as:
@@ III-A3 Loss Function
III-B Progressive Training Strategy
@@
As mentioned before, we argued that the under-exploitation of Transformer-based LTSF models attributed to the heightened similarity within the training samples stemming from rolling forecasting settings. Therefore, we decided to break these similarities during the training process but kept the rolling forecasting settings. Because it does augment the training samples effectively and efficiently. Since increasing stride size for the rolling window does not help with the overfitting issue, inspired by the masking strategy [34], we decided to mitigate the data redundancy issue by adding sample variety. These noises should be added in a manner such that the rate is slower at the beginning to ensure that the model has a good grasp of the data pattern before we start to increase the rate of introducing noises to prevent the model from overfitting. To achieve this goal, we decided to alter dropout rates, as there is a proven connection between dropout and noises [35, 36, 37]. We dynamically alter the dropout rate during training based on a monotone function proposed by [38], whereby the function rate starts slower and later increases as training proceeds. The dropout rate function is based on:
$\theta_{d}(t)=\min(\theta_{max},1-\theta_{max}-{(1-\theta}_{max})\exp(-\gamma t))$ (5)
where t is the $t^{th}$ number of 100 iterations, $\theta_{d}$ is the dropout rate, $\theta_{max}$ is the maximum dropout rate, and $\gamma$ is a hyper-parameter typically set between 0.001 to 0.01.
III-C Seasonal Memory-driven Forecasting
After adding noise to increase the data variety, it is critical to improve the models’ capability to capture long dependencies and intricate temporal patterns. In this paper, we introduce a seasonal memory-driven decoder to facilitate the forecasting process. Applications of the memory concept have demonstrated superior performances in a variety of sequential tasks [39, 40, 41]. In particular, [42] demonstrated that the memory concept is effective in improving LSTM’s ability on multivariate TSF. Therefore, we argued that this concept should also work on Transformer models under the multivariate forecasting setting.
@@ III-C Seasonal Memory-driven Forecasting
There are various ways of implementing memory into Transformers [43, 44, 45, 46, 47]. Our goal is to enhance the Transformer’s ability on both handling increased data complexity brought and recognizing hidden sequential patterns from highly similar training data. We decided to incorporate a memory unit into the decoder block because the decoder block is highly relevant to the output predictions. To achieve this goal, we proposed a seasonal memory, having a self-attention layer that can help to further analyze the hidden sequential patterns that are missed by the encoder-decoder component by combining information stored in the memory matrix together with the model input. The way to connect the seasonal memory block to the decoder is via a memory-driven conditional layer normalization (MDCLN) layer, the general formula for this connection can be written as:
$\rm{MDCLN(Att(h_{l-1}),M_{t})}$ (6)
@@
III-C1 Memory-driven Conditional Layer Normalization
The Memory-driven Conditional Layer Normalization (MDCLN) is a way to integrate seasonal memory into the decoder, where outputs from the decoder’s self-attention layer undergo layer normalization along with the memory information provided from the seasonal memory block. The reason for using layer normalization instead of batch normalization is that input data are normalized before feeding into models as common practice for time-series forecasting. Hence, instead of normalizing values among batches, we use layer normalization to normalize the features within each batch to give each time feature an equal weight for generating predictions.
Figure 3: Overview of the gate mechanism.
@@ III-C1 Memory-driven Conditional Layer Normalization
$\gamma_{t}=\gamma+\it{f_{l}(M_{t})},$ (7)
$\beta_{t}=\beta+f_{l}(M_{t}).$ (8)
@@
$\rm{MDCLN}(\it{h_{l-1}},\it{M_{t}})=\gamma_{t}\odot\frac{h_{l-1}-\mu}{\sigma}+\beta_{t}+f_{l}).$ (9)
@@
III-C2 Seasonal Memory Matrix
@@
$M_{t}=\sigma\left(G_{f}^{t}\right)\ \odot\ M_{t-1}+\sigma\left(G_{i}^{t}\right)\ \odot\ \bar{M_{t}}\rm{,}$ (11)
@@
III-C3 Gate Mechanism
The gate mechanism decides the proportion of new information from $\bar{M_{t}}$ to incorporate and the proportion of $M_{t-1}$ to discard to produce the final $M_{t}$. As shown in Figure. 3, the gate mechanism has two key components: the input gate $G_{i}^{t}$ and the forget gate $G_{f}^{t}$. The two gates are calculated by balancing the embedded decoder input $X_{feed,dec}^{t-1}$ and the previous memory matrix at $t-1$ ($M_{t-1}$), where we duplicate the $embed(X_{feed,dec}^{t-1})$ to match the dimension of $M_{t-1}$. These two gates have the following formulas:
$G_{i}^{t}=embed(X_{feed,dec}^{t-1})\times W_{i}+\tanh{\left(\it{M_{t-1}}\right)}\times\ \it{U_{i}},$ (12)
$G_{f}^{t}=embed(X_{feed,dec}^{t-1})\times W_{f}+\tanh{\left(\it{M_{t-1}}\right)}\times\ \it{U_{f}}.$ (13)
@@ III-C3 Gate Mechanism
TABLE I: The multivariate results comparison of our method with the SOTA Transformer-based LTSF models (three repetitions). The numbers in bold indicate the better-performing models within each Transformer-based LTSF model pair.
@@
IV Experiments
IV-A Datasets
We conduct extensive experiments on 6 public real-life benchmark datasets. For all datasets, train/valid/test sets are split follow the ratio of 0.7/0.1/0.2 in chronological order. ETT (Electricity Transformer Temperature) [15]. The dataset contains information about six power load features and the target value “oil temperature”. We use the two hourly datasets from the source named as $\rm{Etth_{1}}$ and $\rm{Etth_{2}}$. We use the first 20 months of data; Weather [^1] https://www.ncei.noaa.gov/data/local-climatological-data. The dataset contains hourly data about climate information from around 1,600 locations in USA between 2018 and 2020. The data consist of seven quantitative climatological features with “dry builb temperature” as the target; Air Quality [^2] https://archive.ics.uci.edu/ml/datasets/Beijing+Multi-Site+
Air-Quality+Data
. The dataset consists of eleven quantitative hourly features relevant to air quality with the target value set as “PM10” from 12 national air quality monitoring sites. We use the data from the “Wanshouxigong” station; Traffic [^3] https://archive.ics.uci.edu/ml/datasets/Metro+Interstate+
Traffic+Volume
. The dataset contains the traffic condition data measured in the of between Minneapolis and St Paul, Minnesota, USA. The dataset has five hourly numerical features with the target value of “traffic volume”. Exchange [^4] https://github.com/laiguokun/multivariate-time-series-data/blob/master/exchange_rate. The collection of the daily exchange rates of eight foreign countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore ranging from 1990 to 2016.
IV-B Experimental Details
IV-B1 Baselines
We use four Transformer variants as our baseline with generative-style decoding to generate the output sequences in one step [15]. These models are FEDformer[14], AutoFormer[19], Informer[15] and the vanilla Transformer[4], respectively.In addition, since traditional methods like ARIMA and RNNs-based networks lack competitive performance[12, 14], we only compared our model’s performances with the SOTA Transformer-based model, InParformer [12].
TABLE II: Key parameters for backbones.
Prediction Len. {24,48} {168,336} {720}
$d_{model}$ 1024 1024 1024
$d_{ff}$ 2048 2048 2048
n_heads 8 8 8
Dropout 0.1 0.1 0.1
Activation GELU GELU GELU
Batch Size 32 8 4
Enc.Layer no. 1 2 2
Dec.Layer no. 1 1 1
TABLE III: Key parameters for the seasonal memory unit.
AutoFormer Fedformer Informer
$d_{rm}$ 1024 1024 1024
Mem.slot.no. 1 1 1
Mem.head.no. 2 2 4
IV-B2 Implementation Details
IV-B3 Model Reproducibility
To ensure model reproducibility, we present detailed parameter settings in this section. Table II describes the details for all Transformer-based models. We use the same setting for all three Transformer-based models in our experiment because these three models share a very similar structure and show optimal performance under similar settings. However, we use different settings for different prediction output lengths to adjust the model for different prediction tasks. Table II specifies the model dimension ($d_{model}$), the dimension of positional feedforward layers inside the encoder and decoder block ($d_{ff}$), and the number of attention heads in each attention layer (n_heads). In addition, we include information about batch size, layer activation function, the encoder layer number (Enc.Layer no.), and the decoder layer number (Dec.Layer no.). Table III describes the parameters for the relational memory block. There are three key parameters for the relational memory unit: the unit dimension ($d_{rm}$), the number of memory units (Mem.slot.no), and the number of attention heads in the relational memory (Mem.head.no). Here, we apply different relational memory settings for the three Transformer-based models as they require different settings to achieve optimal performance.
We use two evaluation metrics to measure our model performance: Mean Squared Error (MSE) and Mean Absolute Error (MAE).
IV-C Multivariate Results
@@
Table I summarizes the multivariate results of the Transformer-based LTSF models and our proposed methods. We applied the same setting for each transformer-based model pair to achieve a fair comparison. Based on the results in Table I, we find that: (1) Our model can significantly improve the Transformer-based LTSF models’ performances up to 30% on multivariate LTSF problems across most cases; it improves the MSE and MAE scores in 20 of 25 prediction tasks in FEDformer, 23 of 25 prediction tasks in Informer and 19 of 25 tasks in vanilla Transformer. Our results show that our proposed CLMFormer with the progressive training schedule and memory-driven decoder can help Transformers with multivariate LTSF problems. (2) Our method works particularly well in longer prediction settings. For instance, when the prediction length is 720, our approach improves the Informer’s MSE score by 21.63% for the Etth1 dataset and 20.00% for the traffic dataset, in contrast to the 10.00% and 14.96% improvements, respectively, for the prediction length of 24, and this trend is consistently observed across all different models and datasets. The improved performance with longer prediction settings makes sense, considering that the seasonal memory unit is highly effective at dealing with highly similar data, and our progressive training strategy effectively addresses the overfitting issues caused by increased prediction lengths. (3) Our method can further improve the performances of the SOTA LTSF model, FEDformer, on almost all tasks. Such experimental results prove that our method can be seamlessly plugged into varying Transformer-based LTSF systems even when these models are designed to explore the long dependencies and seasonal-trend features to improve the prediction capabilities. The frequency modules in FEDformer can be considered as another kind of memory-driven paradigm, and it is encouraging to observe that our methods bring out additional improvements.
@@
TABLE IV: The univariate results comparison of our method with the Transformer-based LTSF models. The numbers in bold and red indicate the better-performing models within each Transformer-based LTSF model pair and comparing to the InParformer.
@@
TABLE V: Ablation study for the decomposition of our approach, where p.t. is the progressive training scheme and mem refers to our memory approach (three repetitions).
@@
Figure 4: Samples of time series data from the ETTh1 and Air Quality datasets, along with predictions by FedFormer and our CLMFormer. We employ multivariate settings with prediction lengths of 720 and visualize the OT value and temperature value for the ETTh1 and Air Quality datasets, respectively.
IV-D Univariate Results
In Table IV, an extensive univariate results comparison is presented, comparing our method with various Transformer-based LTSF models. Notably, standout performers are indicated by values displayed in bold and red within each Transformer-based LTSF model pair, as well as in contrast to the InParformer model. Since InParformer is not an open-source work, we directly listed the numbers from the original paper[12] for reference. The observations yield several important insights: (1) Our method consistently outperforms alternative Transformer-based models, including FEDformer and AutoFormer, across a spectrum of prediction lengths and datasets, especially in lengthier predictions. This is evidenced by reductions in both MSE and MAE. For instance, in the “Exchange” dataset with a prediction length of 720, our approach built upon FEDformer achieves an MSE of 1.126 compared to FEDformer’s 1.301, while InParformer records 1.172. (2) Our methodology demonstrates pronounced efficacy in extended prediction scenarios, yielding substantial performance gains relative to longer prediction lengths. This pattern is consistently observed across multiple models and datasets. For instance, on the ”ETTh2” dataset with a prediction length of 720, our approach improves an MSE of 0.267 in comparison to FEDformer’s 0.278, and an MAE of 0.398 in comparison to AutoFormer’s 0.409, which also achieves competitive performances to InParformer. (3) The integration of our approach produces enhancements when embedded within the state-of-the-art LTSF model FEDformer and AutoFormer for univariate LTSF, thereby affirming its capability to augment performance in various Transformer-based systems designed to capture extended dependencies and seasonal-trend features. This incorporation leads to additional improvements, complementing existing mechanisms to capture intricate temporal patterns and long dependencies in these models. It is also encouraging to observe the compelling superiority of our proposed methodology in addressing univariate LTSF tasks across diverse contexts.
TABLE VI: Ablation study of different memory matrix initialization methods (three repetitions).

### v5

@@ 3.1.1. Input Embedding
The input data are first passed through an embedding. Different from other types of sequential data, in time series problems, the model should capture the time information. This is extremely crucial in LSTF settings, as models are required to make predictions up to weeks in advance. We decided to utilize the time-series-specific embedding method proposed by (Zhou2020informer, ). The method applies a combination of the context vector, the positional embedding, and the seasonal embedding. The context vector $u^{t}\in R^{L_{S}*d_{model}}$ is obtained by projecting the model input $X^{t}\in R^{L_{S}*d_{f}}$ into a dimension of $d_{model}$ by an 1-D convolutional layer. The goal of the context vector is to transform the input data into an appropriate dimension. Positional embedding (PE(pos)) and seasonal embedding (SE(pos)) are used to capture the local context and global time features. The formula for the input embedding is:
(2) $X_{embed}^{t}=\delta u^{t}+\rm{PE}(\rm{pos})+\sum_{k}[\rm{SE(pos)}]_{k}$
@@
Figure 2. An overview of the model architecture. $M_{t-1}$ represents the memory matrix from the previous prediction, $M_{t}$ represents the updated memory matrix during the current sequence prediction, which in turn is recycled to be the $M_{t-1}$ for the next prediction.
3.1.2. Encoder-Decoder Architecture
The backbone of our method lies in Transformer-based models following an encoder-decoder structure. Despite some minor differences between different variants, their general purpose is the same. The encoder takes past input sequences into the model and generates hidden states as outputs that contain time pattern information, which the decoder uses to make predictions. We decided to use generative-style decoding (Zhou2020informer, ), which has been proven to achieve superior performances in LSTF problems and have exponentially better computational efficiency than conventional dynamic decoding (devlin2018bert, ). The generative-style decoding generates all predictions in one step, with the decoder feeding vector $X^{t}_{feed,dec}$ presented as:
@@ 3.1.3. Loss Function
3.2. Progressive Training Strategy
@@
As mentioned before, we argued traditional Transformer-based models, the strong inductive bias enables them to effectively capture complex temporal dependencies. However, this also means that these models require large-scale datasets for training to generalize well. Unfortunately, most existing TSF benchmarks contain relatively simple and repetitive patterns, which leads to fast convergence and severe overfitting in these Transformer-based models. Therefore, we decided to break these similarities during the training process but kept the rolling forecasting settings. Because it does augment the training samples effectively and efficiently. Since increasing stride size for the rolling window does not help with the overfitting issue, inspired by the masking strategy (devlin2018bert, ), we decided to mitigate the data redundancy issue by adding sample variety. These noises should be added in a manner such that the rate is slower at the beginning to ensure that the model has a good grasp of the data pattern before we start to increase the rate of introducing noises to prevent the model from overfitting. To achieve this goal, we decided to alter dropout rates, as there is a proven connection between dropout and noises (Bishop1995TrainingWN, ; Wager2013DropoutTA, ; Zhai2015DropoutTO, ). We dynamically alter the dropout rate during training based on a monotone function proposed by (Morerio2017CurriculumD, ), whereby the function rate starts slower and later increases as training proceeds. The dropout rate function is based on:
(5) $\theta_{d}(t)=\min(\theta_{max},1-\theta_{max}-{(1-\theta}_{max})\exp(-\gamma t))$
where t is the $t^{th}$ number of 100 iterations, $\theta_{d}$ is the dropout rate, $\theta_{max}$ is the maximum dropout rate, and $\gamma$ is a hyper-parameter typically set between 0.001 to 0.01. Compared with other scheduling methods such as linear or cosine because it offers a gradual and adaptive transition in the dropout rate. The linear scheduling increases the dropout rate at a constant pace, which can lead to a too-aggressive or too-slow regularization effect at different training stages. Cosine scheduling is commonly used for learning rate decay but is less effective for dropout scheduling, as it introduces oscillations that may interfere with stable training progression.
Figure 3. The interaction between curriculum learning and memory-driven mechanism.
3.3. Seasonal Memory-driven Forecasting
We illustrate the interaction between curriculum learning and our memory-driven mechanism in Figure.3. While injecting noises helps mitigate overfitting, it also introduces a new challenge—without proper inductive biases, the model may misinterpret noise as meaningful information, leading to instability in predictions. To counteract this, we introduce a memory-driven mechanism, which serves to reinforce the model’s intrinsic inductive bias. This module ensures that while the model is exposed to progressively noisier training samples, it does not become overly sensitive to noisy data points. Applications of the memory concept have demonstrated superior performances in a variety of sequential tasks (vqamemory2018, ; Ma2020MemoryAG, ; Fei21a, ). In particular, (Xu2020TensorizedLW, ) demonstrated that the memory concept is effective in improving LSTM’s ability on multivariate TSF. Therefore, we argued that this concept should also work on Transformer models under the multivariate forecasting setting.
@@ 3.3. Seasonal Memory-driven Forecasting
There are various ways of implementing memory into Transformers (jiang2019web, ; BaninoBKCZHBBKB20, ; chen2020raio, ; memoryma2021, ; DBLP:conf/nips/ZhouMWW0YY022, ). Our goal is to improve the Transformer’s ability to handle the increased data complexity that is brought about and to recognize hidden sequential patterns from highly similar training data. We decided to incorporate a memory unit into the decoder block because the decoder block is highly relevant to the output predictions. To achieve this goal, we proposed a seasonal memory, having a self-attention layer that can help to further analyze the hidden sequential patterns that are missed by the encoder-decoder component by combining information stored in the memory matrix together with the model input. The way to connect the seasonal memory block to the decoder is via a memory-driven conditional layer normalization (MDCLN) layer, the general formula for this connection can be written as:
(6) $\rm{MDCLN(Att(h_{l-1}),M_{t})}$
@@
3.3.1. Memory-driven Conditional Layer Normalization
The Memory-driven Conditional Layer Normalization (MDCLN) is a way to integrate seasonal memory into the decoder, where outputs from the decoder’s self-attention layer undergo layer normalization along with the memory information provided from the seasonal memory block. We present its detailed architecture in Figure.4 right. The reason for using layer normalization instead of batch normalization is that input data are normalized before feeding into models as common practice for time-series forecasting. Hence, instead of normalizing values among batches, we use layer normalization to normalize the features within each batch to give each time feature an equal weight for generating predictions.
Figure 4. Detailed architectures of the memory-driven conditional layer normalization layer (left, green box) and the gate mechanism (right, purple box).
@@ 3.3.1. Memory-driven Conditional Layer Normalization
(7) $\gamma_{t}=\gamma+\it{f_{l}(M_{t})},$
(8) $\beta_{t}=\beta+f_{l}(M_{t}).$
@@
(9) $\rm{MDCLN}(\it{h_{l-1}},\it{M_{t}})=\gamma_{t}\odot\frac{h_{l-1}-\mu}{\sigma}+\beta_{t}+f_{l}).$
@@
3.3.2. Seasonal Memory Matrix
@@
(11) $M_{t}=\sigma\left(G_{f}^{t}\right)\ \odot\ M_{t-1}+\sigma\left(G_{i}^{t}\right)\ \odot\ \bar{M_{t}}\rm{,}$
@@
3.3.3. Gate Mechanism
The gate mechanism decides the proportion of new information from $\bar{M_{t}}$ to incorporate and the proportion of $M_{t-1}$ to discard to produce the final $M_{t}$. As shown in Figure.4 left, the gate mechanism has two key components: the input gate $G_{i}^{t}$ and the forget gate $G_{f}^{t}$. The two gates are calculated by balancing the embedded decoder input $X_{feed,dec}^{t-1}$ and the previous memory matrix at $t-1$ ($M_{t-1}$), where we duplicate the $embed(X_{feed,dec}^{t-1})$ to match the dimension of $M_{t-1}$. These two gates have the following formulas:
(12) $G_{i}^{t}=embed(X_{feed,dec}^{t-1})\times W_{i}+\tanh{\left(\it{M_{t-1}}\right)}\times\ \it{U_{i}},$
(13) $G_{f}^{t}=embed(X_{feed,dec}^{t-1})\times W_{f}+\tanh{\left(\it{M_{t-1}}\right)}\times\ \it{U_{f}}.$
@@ 3.3.3. Gate Mechanism
4. Experiments
4.1. Datasets
We conduct extensive experiments on 6 public real-life benchmark datasets. For all datasets, train/valid/test sets are split following the ratio of 0.7/0.1/0.2 in chronological order. ETTh (Electricity Transformer Temperature) (Zhou2020informer, ). The dataset contains information about six power load features and the target value “oil temperature”. We use the two hourly datasets from the source named $\rm{Etth_{1}}$ and $\rm{Etth_{2}}$. We use the first 20 months of data, which consists of approximately 17,400 hourly samples per dataset. The six features include electrical load-related attributes such as power factor, active power, and reactive power; Weather [^1] https://www.ncei.noaa.gov/data/local-climatological-data. The dataset contains hourly data about climate information from around 1,600 locations in the USA between 2018 and 2020. The data consist of seven quantitative climatological features with “dry bulb temperature” as the target. Each location contributes approximately 26,300 samples, and the features include variables such as wind speed, humidity, and atmospheric pressure; Air Quality [^2] https://archive.ics.uci.edu/ml/datasets/Beijing+Multi-Site+
Air-Quality+Data
. The dataset consists of eleven quantitative hourly features relevant to air quality with the target value set as “PM10” from 12 national air quality monitoring sites. We use the data from the “Wanshouxigong” station. This dataset contains approximately 35,000 hourly samples, and features include pollutants such as NO2, CO, SO2, and ozone concentrations, in addition to meteorological factors like temperature and wind speed; Traffic [^3] https://archive.ics.uci.edu/ml/datasets/Metro+Interstate+
Traffic+Volume
. The dataset contains the traffic condition data measured between Minneapolis and St Paul, Minnesota, USA. The dataset has five hourly numerical features with the target value of “traffic volume". It consists of around 48,000 hourly samples, and features include traffic speed, lane occupancy, and weather-related variables; Exchange [^4] https://github.com/laiguokun/multivariate-time-series-data/blob/master/exchange_rate. The collection of the daily exchange rates of eight foreign countries including Australia, British, Canada, Switzerland, China, Japan, New Zealand and Singapore ranging from 1990 to 2016. This dataset comprises approximately 9,500 daily samples, with each feature representing the exchange rate of a different currency against the US dollar.
4.2. Experimental Details
4.2.1. Baselines
We use four Transformer variants as our baseline with generative-style decoding to generate the output sequences in one step (Zhou2020informer, ). These models are FEDformer(zhou2022fedformer, ), AutoFormer(Wu2021autoformer, ), Informer(Zhou2020informer, ) and the vanilla Transformer(vaswani2017attention, ), respectively.In addition, since traditional methods like ARIMA and RNNs-based networks lack competitive performance(cao2023inparformer, ; zhou2022fedformer, ), we only compared our model’s performances with the SOTA Transformer-based model, InParformer (cao2023inparformer, ).
Table 1. Key parameters for backbones.
Prediction Len. {24,48} {168,336} {720}
$d_{model}$ 1024 1024 1024
$d_{ff}$ 2048 2048 2048
n_heads 8 8 8
Dropout 0.1 0.1 0.1
Activation GELU GELU GELU
Batch Size 32 8 4
Enc.Layer no. 1 2 2
Dec.Layer no. 1 1 1
Table 2. Key parameters for the seasonal memory unit.
AutoFormer Fedformer Informer
$d_{rm}$ 1024 1024 1024
Mem.slot.no. 1 1 1
Mem.head.no. 2 2 4
4.2.2. Implementation Details
4.2.3. Model Reproducibility
To ensure model reproducibility, we present detailed parameter settings in this section. Table.1 describes the details for all Transformer-based models. We use the same setting for all three Transformer-based models in our experiment because these three models share a very similar structure and show optimal performance under similar settings. However, we use different settings for different prediction output lengths to adjust the model for different prediction tasks. Table.1 specifies the model dimension ($d_{model}$), the dimension of positional feedforward layers inside the encoder and decoder block ($d_{ff}$), and the number of attention heads in each attention layer (n_heads). In addition, we include information about batch size, layer activation function, the encoder layer number (Enc. Layer no.), and the decoder layer number (Dec.Layer no.). Table.2 describes the parameters for the relational memory block. There are three key parameters for the relational memory unit: the unit dimension ($d_{rm}$), the number of memory units (Mem.slot.no), and the number of attention heads in the relational memory (Mem.head.no). Here, we apply different relational memory settings for the three Transformer-based models as they require different settings to achieve optimal performance.
We use two evaluation metrics to measure our model performance: Mean Squared Error (MSE) and Mean Absolute Error (MAE). MSE and MAE are among the most commonly used metrics for TSF due to their ability to measure prediction accuracy effectively. MAE measures the absolute difference between predicted and actual values. It is robust to outliers, as all errors contribute equally to the final value. MSE penalizes larger errors more heavily, making it sensitive to large deviations. It ensures that the model prioritizes reducing larger errors, which is important in forecasting tasks where extreme deviations can be critical. We conduct paired t-tests to compare our model’s performance against baselines. The p-values from these tests help determine whether the observed improvements are statistically significant. To clearly indicate statistical significance in our results, we use (*) in our tables to denote significant differences ($p<0.01$).
@@ 4.2.3. Model Reproducibility
Table 3. The multivariate results comparison of our method with the SOTA Transformer-based LTSF models (three repetitions). The numbers in bold indicate the better-performing models within each Transformer-based LTSF model pair.
@@
4.3. Multivariate Results
@@
Table.3 summarizes the multivariate results of the Transformer-based LTSF models and our proposed methods. We applied the same setting for each transformer-based model pair to achieve a fair comparison. Based on the results in Table.3, we find that: (1) Our model can significantly improve the Transformer-based LTSF models’ performances up to 30% on multivariate LTSF problems across most cases; it improves the MSE and MAE scores in 20 of 25 prediction tasks in FEDformer, 23 of 25 prediction tasks in Informer and 19 of 25 tasks in vanilla Transformer. Our results show that our proposed CLMFormer with the progressive training schedule and memory-driven decoder can help Transformers with multivariate LTSF problems. (2) Our method works particularly well in longer prediction settings. For instance, when the prediction length is 720, our approach improves the Informer’s MSE score by 21.63% for the Etth1 dataset and 20.00% for the traffic dataset, in contrast to the 10.00% and 14.96% improvements, respectively, for the prediction length of 24, and this trend is consistently observed across all different models and datasets. The improved performance with longer prediction settings makes sense, considering that the seasonal memory unit is highly effective at dealing with highly similar data, and our progressive training strategy effectively addresses the overfitting issues caused by increased prediction lengths. (3) Our method can further improve the performances of the SOTA LTSF model, FEDformer, on almost all tasks. Such experimental results prove that our method can be seamlessly plugged into varying Transformer-based LTSF systems even when these models are designed to explore the long dependencies and seasonal-trend features to improve the prediction capabilities. The frequency modules in FEDformer can be considered as another kind of memory-driven paradigm, and it is encouraging to observe that our methods bring out additional improvements.
@@
Table 4. The univariate results comparison of our method with the Transformer-based LTSF models. The numbers in bold and red indicate the better-performing models within each Transformer-based LTSF model pair and compared to the InParformer. These improvements are statistically significant at $p=0.05$.
@@
Table 5. Ablation study for the decomposition of our approach, where p.t. is the progressive training scheme and mem refers to our memory approach (three repetitions).
@@
4.4. Univariate Results
In Table.4, an extensive univariate results comparison is presented, comparing our method with various Transformer-based LTSF models. Notably, standout performers are indicated by values displayed in bold and red within each Transformer-based LTSF model pair, as well as in contrast to the InParformer model. Since InParformer is not an open-source work, we directly listed the numbers from the original paper(cao2023inparformer, ) for reference. The observations yield several important insights: (1) Our method consistently outperforms alternative Transformer-based models, including FEDformer and AutoFormer, across a spectrum of prediction lengths and datasets, especially in lengthier predictions. This is evidenced by reductions in both MSE and MAE. For instance, in the “Exchange" dataset with a prediction length of 720, our approach built upon FEDformer achieves an MSE of 1.126 compared to FEDformer’s 1.301, while InParformer records 1.172. (2) Our methodology demonstrates pronounced efficacy in extended prediction scenarios, yielding substantial performance gains relative to longer prediction lengths. This pattern is consistently observed across multiple models and datasets. For instance, on the "ETTh2" dataset with a prediction length of 720, our approach improves an MSE of 0.267 in comparison to FEDformer’s 0.278, and an MAE of 0.398 in comparison to AutoFormer’s 0.409, which also achieves competitive performances to InParformer. (3) The integration of our approach produces enhancements when embedded within the state-of-the-art LTSF model FEDformer and AutoFormer for univariate LTSF, thereby affirming its capability to augment performance in various Transformer-based systems designed to capture extended dependencies and seasonal-trend features. This incorporation leads to additional improvements, complementing existing mechanisms to capture intricate temporal patterns and long dependencies in these models. It is also encouraging to observe the compelling superiority of our proposed methodology in addressing univariate LTSF tasks across diverse contexts.
Table 6. Ablation study of different memory matrix initialization methods (three repetitions).